22

a. On the basis of the three individual demand schedules below, and assuming these three people are the only ones in the society, determine the collective demand schedule on the assumption that the good is a public good. Instructions: Enter your answers as whole numbers. Demand Individual #1 Individual #2 Individual #3 Public Good Price Qd Price Qd Price Qd Price Qd $8 S8 S8 2 4 4 4 4 4 2b. Use the public demand schedule above and the following supply schedule to ascertain the optimal quantity of this public good. Quantity Supplied 10 Price S25 24 23 21 20 0 Optimal quantity
Answer #1

a.

Public Good: Price Quantity, Qd
8 + 8 + 8 = 24 1
7 + 7 + 7 = 21 2
6 + 6 + 6 = 18 3
5 + 5 + 5 = 15 4
4 + 4 + 4 = 12 5
3 + 3 + 3 = 9 6
2 + 2 + 2 = 6 7
1 + 1 + 1 = 3 8

(Demand for public goods is vertical summation of individual demand curves which means price by each individual is added for every quantity.)

b. Optimal quantity = 2
(As Qd = quantity supplied = 2 at price 21)